OpenStudy (anonymous):

f (x) = x−4/5 − 5, [−32, 32] is f continous and differentiable on the closed interval?

OpenStudy (anonymous):

or is it not

OpenStudy (amistre64):

reposting a question that makes no sense to begin with is not gong to make it more sensible :)

OpenStudy (amistre64):

if we take the notation "as is" then this is just f(x) = x + C, which is differentiable and continuous

OpenStudy (anonymous):

is the function continous and differtiable

OpenStudy (amistre64):

yes it is and no it isnt ....

OpenStudy (anonymous):

Dove, please make correct use of parens to disambiguate your notation.

OpenStudy (amistre64):

\[f(x) = x−\frac{4}{5} − 5\] yes \[f(x) = \frac{x−4}{5 − 5}\] no \[f(x) = \frac{x−4}{5} − 5\] yes \[f(x) = x−\frac{4}{5 − 5}\] no